m. tropine is available in a 400 μg/ml solution. To give 1.2 mg, how many ml must you give?
Answer #1

ANSWER

   Amount of (ml) drug need to be administered.

       Formula.

        Amount of drug administered = Amount desired (D) / Amount in hand (H) * Quantity (Q).

              Here

                       Amount desired = 1.2 mg.

                     Amount in hand = 400 mcg.( 0.4 mg).

                           Quantity = 1 ml.

                   So,

                     Amount of atropine administerd = D / H *Q

                                                                          = 1.2 / 0.4 * 1

                                                                          = 3 * 1

                                                                          = 3 ml.( 1.2 mg)